OpenStudy (anonymous):

[2.03] MA.7.A.3.4 Match each equation with its equivalent equation. Match Term Definition six x equals twentyfour A) four x equals eight x divided by six equals six B) 4x=16 x divided by two equals sixteen C) x divided by twelve equals three eight x equals sixteen D) x divided by eight equals four

OpenStudy (anonymous):

@BrAnDoN_AAS

OpenStudy (anonymous):

what grade is this math google links arent being usefull for this thing

OpenStudy (anonymous):

7th

OpenStudy (anonymous):

oh haha im in seventh

OpenStudy (anonymous):

cool

OpenStudy (anonymous):

1. C 4. A

OpenStudy (anonymous):

well first you might want to convert it all into math speak >,< :P 6x = 24 A. x/12 = 3 x/6 = 6 B. 4x = 16 x/2 = 16 C. x/8 = 4 8x = 16 D. 4x = 8

OpenStudy (anonymous):

@sangya21

OpenStudy (anonymous):

to solve for x for 6x = 24 divide both sides by 6 so x = 24/6 simplified x/2 = 16 multiply both sides by 2 x = 16*2 x/2 = 16 multiply both sides by 2 8x = 16 divide both sides by 8

OpenStudy (anonymous):

the rest

OpenStudy (anonymous):

ake for example the first one: 6x = 24 in order to find another one that is the same as this equation what is needed to multiply 6 * ___ = 24

OpenStudy (anonymous):

the answer is 4, so now look at the equations on the right which one gives you a correct value if you put a 4 for x

OpenStudy (anonymous):

hat would be 4x = 16 because 4*4 = 16 since the equations match their x value should be the same and will give you a correct answer

OpenStudy (anonymous):

so the first one goes with B

OpenStudy (anonymous):

THE SECOND ONE GOES WITH A

OpenStudy (anonymous):

THE NEXT ONE GOES WITH d

OpenStudy (anonymous):

copying and pasting are we now brandon

OpenStudy (anonymous):

3rd is c and 4th is D

OpenStudy (anonymous):

your welcome

OpenStudy (anonymous):

XD